# Setting the Course, but Forgetting to Steer: Analyzing Compliance with GDPR’s Right of Access to Data by Instagram, TikTok, and YouTube

### v1

@@ Abstract
The comprehensibility and reliability of data download packages (DDPs) provided under the General Data Protection Regulation’s (GDPR) right of access are vital for both individuals and researchers. These DDPs enable users to understand and control their personal data, yet issues like complexity and incomplete information often limit their utility. Also, despite their growing use in research to study emerging online phenomena, little attention has been given to systematically assessing the reliability and comprehensibility of DDPs.
To bridge this research gap, in this work, we perform a comparative analysis to assess the comprehensibility and reliability of DDPs provided by three major social media platforms, namely, TikTok, Instagram, and YouTube. By recruiting 400 participants across four countries, we assess the comprehensibility of DDPs across various requirements, including conciseness, transparency, intelligibility, and clear and plain language. Also, by leveraging automated bots and user-donated DDPs, we evaluate the reliability of DDPs across the three platforms. Among other things, we find notable differences across the three platforms in the data categories included in DDPs, inconsistencies in adherence to the GDPR requirements, and gaps in the reliability of the DDPs across platforms. Finally, using large language models, we demonstrate the feasibility of easily providing more comprehensible DDPs.
  *These authors contributed equally to this work.
@@ 1 Introduction
Ensuring compliance with the General Data Protection Regulation (GDPR) [12] is essential for platforms handling personal data, as it establishes legal requirements for data processing, security, and user rights. One of these rights is the right of access by data subjects, which enlists provisions whereby data subjects (i.e., end users) can request access to their data being processed by platforms (Article 15(3) GDPR). Platforms have to comply with these requests within a stipulated period of one month and provide the user with a copy of their data. Usually, platforms provide this data in the form of data download packages (DDPs) [34]. These DDPs often include information about the users’ activities on and off the platform, device and app information, inferred preferences, and other information depending on the business of the platform.
Implementation across platforms: In fact, the GDPR stipulates that information shall be provided to users in commonly used electronic forms (Art. 15(3)(2) GDPR). Such an implementation is crucial for enabling users to exercise their right to data portability (Article 20), which allows end users to transmit data from one platform to the other without hindrance [12]. However, the GDPR does not provide explicit standards for the structure or content of DDPs, leading to varying implementations of Article 15 across platforms [8].
@@
Understandably, platforms that offer different types of services interpret and implement the GDPR provisions differently. For instance, the DDP of a digital marketplace like Amazon is naturally distinct from that of a digital public space like TikTok. However, platforms offering similar services should be able to provide comparable information to users. For example, platforms providing short-format video streaming services should, in principle, offer similar details through their DDPs. To the best of our knowledge, such a comparative study of the implementation of the right of access under the GDPR is rare in the literature. This observation leads to a key unexplored research question as follows: RQ1: How similarly do platforms offering similar services implement the right of access under the GDPR (Article 15(3))?
@@
Desideratum 1: Comprehensibility: Irrespective of the answer to RQ1, these implementations (DDPs) should adhere to a set of requirements outlined in the GDPR to ensure user comprehensibility. According to Article 12(1), these DDPs should be (a) concise, (b) transparent, (c) intelligible, (d) easily accessible, and (e) using clear and plain language [12]. However, neither the GDPR nor accompanying guidelines [14] provide a clear technical definition of these criteria. Further, adherence to these requirements in the current implementations also remains unexplored in the research community. The authors in [8] came closest to understanding the question of adherence, but they restricted their analyses to conciseness. This gap leads us to the second research question in this paper: RQ2: How comprehensible are the DDPs provided by platforms under Articles 12 and 15 of the GDPR?
Desideratum 2: Reliability: Finally, another important expectation for the effective implementation of the provision is that the information in the DDPs is reliable. With the increased restriction on access to platform data for research [33, 4], many researchers are relying on data elicited through exercising the right of access for understanding the dynamic interplay between population and platforms. In these so-called “data donations” based studies, consenting participants first request their own data from the platforms, and then provide this data to researchers, typically for financial compensation. Several recent scientific studies use this setup to explore fundamental questions about digital platforms— such as user engagement [38], content recommendation [35], and ad targeting processes [36] — and even questions beyond online activities, e.g., offline risks [3] or the likelihood of addiction [37] using the resultant DDPs. However, if the reliability of the information provided by the platforms is compromised, then the insights drawn from such data sources might be invalid.
To this end, a recently conducted survey among researchers indicates that 86% of researchers report concerns regarding the reliability of the data collected using DDPs [34]. The most important dimensions of reliability that researchers are concerned about include: (a) completeness, (b) correctness, and (c) consistency of the shared information. This brings us to our final research question: RQ3: How reliable is the information within the DDPs provided by platforms under Article 15(3) of the GDPR?
The current work: Although the current work could have been conducted on any digital platform, due to the growing popularity in consumption of short–format videos [38, 25] and the impact they have on the population [13], we investigate the aforementioned research questions on Instagram, TikTok, and YouTube.  Figure 1 describes the pipeline of our conducted research. After analyzing the DDPs shared by these platforms, we conducted an extensive survey among 400 participants from Germany, France, Italy, and Spain to evaluate their assessment of the comprehensibility of the respective DDPs. Further, by utilizing (a) participants’ and (b) European Data Protection Board’s interpretations of the requirements, (c) current advancements in large language models (LLMs), and (d) our technical expertise, we propose data representations for the different categories of information. These recommended representations are then evaluated against the best representations of respective categories in the current implementation of the platforms by 200 participants.
To answer the questions regarding the correctness and completeness of the DDPs, we browse the three platforms using sock-puppet accounts to (a) automatically log the browsing behavior and (b) request the DDPs from each of the platforms. We repeat this process for a period of one month to understand the consistency in the shared DDPs for the accounts. In addition, to better understand the consistency across various accounts, we also collect DDPs requested by real-world participants from the three platforms.
@@
Figure 1: Our pipeline to evaluate comprehensibility and reliability of the current evaluation of Article 15(3) of the GDPR.
We summarize our major findings with respect to each research question below.
RQ1: The studied platforms implement Article 15 of the GDPR differently: Instagram, TikTok, and Youtube share different categories of data in their DDPs. While Instagram and TikTok share more data with their users, YouTube’s shared data is rather limited. Irrespective of the amount of data, the only underlying similarity is that none of the platforms share the purpose of data collection and processing in their DDPs.
RQ2: Instagram’s current data representation is the most comprehensible: Based on the preference votes of 400 participants, we find Instagram’s DDP adheres to GDPR’s requirements more than the other two platforms.
RQ2: Considering interpretation from different stakeholders significantly improves the data representation: 200 survey participants prefer our proposed data representation to current representations along all requirements except conciseness. In fact, our proposal is preferred in as many as 44 out of the 52 evaluations (13 categories $\times$ 4 requirements).
@@
RQ3: TikTok’s DDP is the most reliable as per our evaluation: Our analyses using sock-puppet accounts reveal TikTok’s DDP presents a complete, correct, and consistent representation of our browsing behavior. Instagram and YouTube exhibit varying degrees of missing data across DDPs collected over the period of our study. However, the analyses among real users show Instagram and TikTok share data for different periods in their DDPs for different categories, whereas YouTube shares all the data types for similar duration.
To the best of our knowledge, this study is the first detailed evaluation of adherence of different requirements mentioned in the GDPR. The insights drawn from our study further underline the need for standardization and follow-up on the implementations to ensure that the primary goals of data protection laws are achieved.
@@ 2 Background and Related Work
To position the current work in the perspective of the literature, we present a brief overview of the following: (a) GDPR and the rights of access, (b) leveraging data download packages for platform audits, (c) audits of implementations of provisions of the GDPR.
@@
GDPR and the right of access: The GDPR is a landmark privacy law designed to give individuals greater control over their personal data and to standardize data protection rules across the European Union (EU) [12]. While the GDPR has many provisions and rights enlisted, our work focuses on Article 15(3): the right of access, which allows users to request a copy of their data collected by online platforms. Most platforms implement this requirement by providing a Data Download Package (DDP) upon request. For such implementations, GDPR provides a set of desiderata for the DDPs. Specifically, Article 12(1) requires a DDP to be concise, intelligible, transparent, accessible, and in clear and plain language [12]. Moreover, the GDPR gives the right to end-users to be able to port their data from one platform to another under Article 20. However, if platforms offering similar services implement these rights in significantly different ways, platforms will essentially deprive end users of their right to portability.
@@
Leveraging data download packages: The richness of DDPs has enabled new research directions previously deemed infeasible. Indeed, many recent studies are utilizing them through donations made from participants with their explicit consent. To this extent, researchers have used DDPs to conduct research on topics related to personal health and safety [20, 37, 3], news and politics [6, 17, 18], auditing recommendation algorithms [24, 35], analyzing user behavior on social media platforms [38, 15], and understanding ad targeting types and user perceptions [36].
Audits of implementations of GDPR: Many studies have investigated (a) whether consent forms are designed properly by platforms [29, 30, 23, 9], (b) whether platforms are undermining user privacy under the pretense of ‘legitimate interest’ [31, 21, 22], (c) inconsistencies in platform’s data withdrawal behavior [11]. However, little attention has been given to platforms’ implementation of the right to access. A recent study shows that end users often view DDPs as overwhelming and unclear, leaving key questions unanswered while triggering privacy concerns [8]. Moreover, researchers have expressed concerns about the reliability of the DDPs and have faced challenges in using them [34, 19].
The current work: To the best of our knowledge, our work is one of the first studies that tries to unravel the current implementations of Article 15(3) of the GDPR from the lens of comprehensibility and reliability. Furthermore, our proposed recommendations for the potential implementation of DDPs may not only improve the current implementations, but also help in the implementation of other rights provided to users under the GDPR, such as the right of data portability.
3 RQ1: How similarly do platforms offering similar services implement GDPR Article 15?
@@
Before analyzing the comprehensibility and reliability of DDPs, we first look into the content of the DDPs provided by each of the platforms (i.e., we compare how the platforms are currently implementing GDPR Article 15(3)).
@@ 3.1 Platforms under consideration
Since the primary focus of the study is to understand how similar platforms implement the same provision in the GDPR, we decided to focus on multiple platforms providing similar services. To this end, there are many options for digital streaming platforms (e.g., Netflix, Disney, etc.), digital marketplaces (e.g., Amazon, Zalando, etc.), and social networking platforms (e.g., Facebook, X, LinkedIn, etc.). In this study, we focus on three platforms that provide short-format video streaming services: TikTok, Instagram, and YouTube. Our motivation to study these platforms stems from the recent increase in popularity of short-format video platforms [38, 25]. Furthermore, all three platforms have been designated as Very Large Online Platforms (VLOPs) by the EU under the recently enacted Digital Services Act (DSA), indicating the far-reaching impact that these platforms have on the European society [13].
TikTok, Instagram, and YouTube implement Article 15(3) of the GDPR, through which they share DDPs with their end-users upon request. In the current implementations, platforms share data primarily in two formats (a) machine-readable – which is usually a single JSON or a collection of JSON files; (b) human-readable – where Instagram and YouTube opt for HTML version, whereas TikTok shares it in TXT files segregated across many directories. If one takes a closer look at the DDPs received from any of the platforms, their contents can be broadly divided into the following categories: (a) user’s usage, (b) user’s content, (c) personal details, (d) advertisements, (e) miscellaneous. Next, we examine the DDPs that Instagram, TikTok, and YouTube provide and observe their content and presentation. Most of the observations discussed below are summarized in Table 5 (see Appendix A).
3.2 Information regarding users’ usage
One of the most important facets of the information contained within the DDPs is how a user uses the social media platform. Such usage information may include the content that a user watches, searches for, or engages with implicitly or explicitly.
Watch history: Watch history is an ordered list of all the content that a user has watched before requesting the data. DDPs of all three platforms contain watch history data of the data subject. However, as Table 1 shows the data provided across the three platforms are widely varying. Among these, two points are strikingly different – (a) while YouTube provides this data for the entire lifetime of a user on the platform, Instagram provides data of at most the last two weeks and TikTok provides data for around six months before the data request (without any clear explanation for the reduced duration); (b) while TikTok and YouTube provide the watch history all at once, Instagram segregates it into three files based on the type of content, i.e., ads, posts, and videos watched by the user.
@@ 3.2 Information regarding users’ usage
Table 1: Comparison of watch history data across platforms.
Like history: Similar to watch history, like history is an ordered list of all the contents that a user has liked before requesting the data. The details across the three platforms are noted in Table 2. The table demonstrates that there is a clear lack of consistency in how platforms currently implement Article 15(3) GDPR. While one platform (YouTube) records the like history on its application, it does not share this information in the YouTube DDP  [^1] Although Like history is not there in YouTube ddp, it is included in the Google DDP.. Another platform (Instagram) provides the watch history for just a couple of weeks without any link to the watched content, but it provides the like history of a lifetime, including references to the liked content.
@@
Table 2: Comparison of like history details across platforms.
Time spent on platform: While all three platforms have a detailed dashboard that mentions the average time spent and the exact amount of time users spend on the application daily, the same information is not provided in the DDP in response to Article 15 GDPR. On a positive note, TikTok recently added a new field in the data, namely “Activity Summary,” which enlists how many videos a user has commented on, shared, or watched until the end since their registration.
Other usage activities on platform:  Table 5 (in Appendix A) shows some other important usage activities. Most of the activities, e.g., comment, search, save, share, and writing a message are recorded by all platforms where the features are applicable. However, sharing across applications, i.e., when a user shares the video on other social media platforms or copies the link for posting elsewhere, is only shared in TikTok DDPs. Also, while both TikTok and Instagram maintain a list of inferred interests of the data subject, YouTube DDPs do not have them.
3.3 Information regarding user’s content

### v2

@@ Abstract
The GDPR’s Right of Access aims to empower users with control over their personal data via Data Download Packages (DDPs). However, their effectiveness is often compromised by inconsistent platform implementations, questionable data reliability, and poor user comprehensibility. This paper conducts a comprehensive audit of DDPs from three social media platforms (TikTok, Instagram, and YouTube) to systematically assess these critical drawbacks. Despite offering similar services, we find that these platforms demonstrate significant inconsistencies in implementing the Right of Access, evident in varying levels of shared data. Critically, the failure to disclose processing purposes, retention periods, and other third-party data recipients serves as a further indicator of non-compliance. Our reliability evaluations, using bots and user-donated data, reveal that while TikTok’s DDPs offer more consistent and complete data, others exhibit notable shortcomings. Similarly, our assessment of comprehensibility, based on surveys with 400 participants, indicates that current DDPs substantially fall short of GDPR’s standards. To improve the comprehensibility, we propose and demonstrate a two-layered approach by: (1) enhancing the data representation itself using stakeholder interpretations; and (2) incorporating a user-friendly extension (Know Your Data) for intuitive data visualization where users can control the level of transparency they prefer. Our findings underscore the need for clearer and non-conflicting regulatory guidance, stricter enforcement, and platform commitment to realize the goal of GDPR’s Right of Access.
The supplementary materials and code are available at https://github.com/saikeerthana00/GDPR_Compliance. This work has been accepted at IEEE S&P 2026 as a full paper.
@@ 1 Introduction
The General Data Protection Regulation (GDPR) [10] establishes legal requirements for data processing, security, and user rights. One of these rights is the Right of Access by data subjects, which includes provisions that allow data subjects (i.e., users) to request access to their data being processed by platforms (Article 15(3) GDPR). Along with a copy of such data, GDPR also requires platforms to communicate the purpose of data collection, period of retention, and other recipients [^1] As per Article 4 (9) of the GDPR, ‘recipient’ means a natural or legal person, public authority, agency or body, to which personal data are disclosed [10]. of such data (Article 15(1) GDPR). Platforms have to comply with these requests within a stipulated period of one month. Platforms usually provide this data in the form of data download packages (DDPs) [33].
@@
Implementation across similar platforms: However, the GDPR does not provide standards for the structure or content of DDPs, leading to varying implementations of Article 15 across platforms [6]. Understandably, platforms that offer different types of services are expected to interpret and implement these GDPR provisions differently. For instance, the DDP of a digital marketplace like Amazon is naturally distinct from that of a digital public space like TikTok. However, platforms offering similar services should be able to provide comparable information to users. For example, platforms providing short-format video streaming services should, in principle, offer similar details in their DDPs. Such consistent implementation across similar platforms is one of the important factors for enabling users to exercise their Right to Data Portability (Article 20 GDPR), which allows users to transmit their data across platforms without hindrance [10]. Therefore, there is a need for systematic and comparative audit of implementations across platforms. To the best of our knowledge, such a comparative study of the implementation of the GDPR’s Right of Access across similar platforms is rare in the literature. These concerns lead to the following key unexplored research question:
RQ1: How similarly do platforms offering similar services implement and thus comply with GDPR’s Right of Access?
Reliability of DDPs’ content: Irrespective of the similarity of implementations, a fundamental concern is the reliability of the content in DDPs. Lack of reliability of the DDPs undermines the overarching purpose of the Right of Access and severely compromises their utility towards users. Furthermore, unreliable DDPs can create a distorted understanding of data processed by platforms, limiting users’ digital autonomy and right to digital privacy. Such lack of reliability can also have ramifications for other stakeholders too. Researchers increasingly rely on these DDPs for studying the dynamic interplay between population and platforms [39, 38, 35, 36]. To this end, incorrect or inconsistent information in platforms’ DDPs risks the insights drawn from such studies being fundamentally flawed and misleading. In fact, the above concern is no longer a mere theoretical one. A recently conducted survey among researchers who have used DDP in their research indicates that 86% of them have reliability concerns with DDPs [33]. Their concerns are primarily centered around the following dimensions – (a) completeness: ensuring all relevant data is present in DDPs for the stipulated period, (b) correctness: verifying the accuracy of the information, such as timestamps and event sequences recorded in DDPs, and (c) consistency: assessing uniformity of DDPs across different requests for the same user or across users in different locations. While researchers have reported these concerns, there is no prior systematic audit of the reliability of the DDPs, leading to our second research question:
RQ2: How reliable is the information within the DDPs provided by platforms?
@@
Comprehensibility of DDPs: Even if DDPs are to contain accurate and complete information, their utility would be severely limited if users cannot easily understand the content presented to them i.e., if DDPs are not comprehensible for end-users. In fact, to ensure comprehensibility of the communication, Article 12(1) of the GDPR states these DDPs should be (a) concise, (b) transparent, (c) intelligible, (d) easily accessible, and (e) using clear and plain language [10]. However, neither the GDPR nor accompanying guidelines [12] provide any clear technical definition of these requirements. Additionally, adherence to these requirements in the current implementations also remains unexplored in the research community. The authors in [6] came closest to understanding the question of adherence, but they restricted their analyses to conciseness. This gap leads to the third research question in this paper:
RQ3: (a) How comprehensible are the DDPs provided by platforms and (b) how can their comprehensibility be improved?
@@
Figure 1: Pipeline to evaluate comprehensibility and reliability of the implementation of Article 15(3) of the GDPR.
The current work: Although the current work could have been conducted on any digital platform, due to the growing popularity of short–format videos [39, 23] and their impact [11], we investigate the discussed research questions on TikTok, Instagram, and YouTube. Figure 1 describes the pipeline of our conducted research.
To audit reliability of DDPs, we browse the three platforms using sock-puppet accounts to (a) automatically browse and log the behavior and (b) request the corresponding DDPs from each of the platforms. We quantify correctness and completeness by comparing our log with the DDPs from each platform. We repeat this process for a period of one month to understand the consistency in the shared DDPs for the accounts. Also, to better understand the consistency across various accounts, we collect DDPs requested by real-world participants from the three platforms.
To assess the comprehensibility of DDPs, we conduct an extensive survey among 400 participants from Germany, France, Italy, and Spain. Beyond understanding the comprehensibility of the DDPs, we also ask participants about their interpretation of different GDPR prescribed requirements. One of the key observations in this survey is that some of the requirements are at odds with each other.
To this end, by utilizing interpretations of participants and European Data Protection Board, we propose data representations for the different categories of information that aim at striking a better balance among these requirements. Furthermore, to improve the comprehensibility of the DDPs, we propose a browser extension (Know Your Data) which (a) enables users to request and download their data in one click; (b) provides them the autonomy to choose their desired level of transparency in data visualization. The code is made publicly available. Below, we summarize our major findings.
RQ1: Instagram, TikTok, and YouTube interpret and implement GDPR’s Right of Access differently, with YouTube offering surprisingly limited data compared to the others. Critically, none adhere to the fundamental requirement of (a) disclosing the purpose of data collection, (b) duration of retention, (c) other entities with whom the data is shared for different data categories within DDPs.
@@
RQ2: Our analyses using sock-puppet accounts reveal TikTok’s DDP presents a complete, correct, and consistent representation of our logged browsing behavior. Instagram and YouTube exhibit varying degrees of missing data across DDPs collected during our study. However, the analyses among real users show Instagram and TikTok share data for different periods in their DDPs for different categories, whereas YouTube shares all data for a similar duration.
RQ3(a): In our user survey, we find Instagram’s DDP to be more comprehensible than that of the others. However, survey participants’ interpretation of the different requirements reveals some of them may have conflicting goals; e.g., while participants interpret conciseness to prioritize relevant information, they interpret transparency to provide complete information with full disclosure.
RQ3(b): Improving comprehensibility requires a layered approach. One can improve the data representation themselves by accounting for interpretations from different stakeholders to achieve a better trade-off between the conflicting requirements. Furthermore, to improve the understandability, an additional layer of visualization in the form of a dashboard may improve the current state-of-the-art by enabling users to decide the level of conciseness/transparency they prefer.
Implications: Our findings reveal several instances of platform-specific and systemic non-compliance, providing regulators with starting points to investigate the platforms. Furthermore, insights drawn from our surveys highlight the conflicting nature of GDPR’s comprehensibility requirements making it difficult for platforms to adhere to all of them at once. At the same time, our recommendations underline the need for standardization and improved DDP communication through dashboards to ensure that the primary goals of data protection laws are achieved.
@@ 2 Background and related work
GDPR and rights of data subjects: The GDPR is a landmark privacy law designed to give individuals greater control over their personal data and to standardize data protection rules across the European Union (EU) [10]. While the GDPR has many provisions and rights enlisted, our work focuses on Article 15: the Right of Access, which allows users to access their data collected and processed by online platforms. Most platforms implement this requirement by providing a Data Download Package (DDP) upon request.
For such implementations, GDPR provides a set of desiderata for the DDPs. Specifically, Article 12(1) requires a DDP to be concise, intelligible, transparent, accessible, and in clear and plain language [10]. Moreover, Article 20 gives the right to end-users to port their data from one platform to another. However, if platforms offering similar services implement these rights in significantly different ways, platforms will essentially deprive end users of their right to portability.
@@
Leveraging DDPs for platform audits: Even at the current level of implementation, the richness of DDPs has enabled new research directions previously deemed infeasible. Many recent studies are utilizing them through donations made from participants with their explicit consent. To this extent, researchers have used DDPs to conduct research on topics related to personal health and safety [18, 38, 3], news and politics [5, 15, 16], auditing recommendation algorithms [22, 35], analyzing user behavior [39, 13], and ad targeting on social media platforms [36].
GDPR compliance audits: Since the enforcement, numerous academic evaluations have been conducted to assess compliance with different provisions of GDPR. Prior works have investigated (a) effective implementation of data portability [37, 32], (b) ethical design of consent forms [28, 29, 21, 7], (c) privacy violations under the pretense of ‘legitimate interest’ [30, 19, 20], and (d) inconsistencies in platform’s data withdrawal behavior [9].
Current work: However, little attention has been given to platforms’ implementation of the right to access. While there are some studies raising privacy [6] and reliability concerns [33, 17], none of the studies systematically audit the Right of Access on the legal requirements proposed by the GDPR itself. To the best of our knowledge, this work is one of the first studies that tries to empirically evaluate the current implementations of Article 15 of the GDPR from the combined lens of comprehensibility and reliability. Besides auditing the compliance, we also propose a number of recommendations for policymakers and platforms alike to improve the comprehensibility for end-users.
3 Auditing compliance of similar platforms
@@
Before analyzing reliability and comprehensibility, we first compare how platforms are currently implementing GDPR Article 15 by looking at the content of their DDPs and asses their compliance with the prescribed requirements.
@@ 3.1 Platforms under consideration
In this study, we focus on three platforms that provide short-format video streaming services: TikTok, Instagram, and YouTube. Our motivation to study these platforms stems from the recent popularity of short-format video platforms [39, 23]. Furthermore, all three platforms have been designated as Very Large Online Platforms (VLOPs) by the European Union (EU) under the recently enacted Digital Services Act (DSA), indicating the far-reaching impact that these platforms have on the European society [11].
What do these platforms share?: Upon users’ request, these platforms share Data Download Packages (DDPs) primarily in two formats (a) machine-readable – a single or a collection of JSON/CSV file(s); (b) human-readable – Instagram and YouTube opt for HTML version, whereas TikTok shares in TXT files segregated across directories.
If one takes a closer look at the DDPs received from any of the platforms, their contents can be broadly divided into the following categories: (a) user’s engagement, (b) personal details, (c) advertisements etc. Next, we examine the DDPs that TikTok, Instagram, and YouTube provide and observe their degree of compliance.
@@
TABLE I: Comparison of watch history across platforms.
@@
TABLE II: Comparison of like history across platforms.
Information Type TikTok Instagram YouTube Minimum expected fields
User’s usage Watch Y N Y Content Id, ts
Search Y Y Y Query term, ts
Comment $N^{*}$ Y Y Comment text, Content Id, ts
Like Y Y $Y^{g}$ Content Id, ts
Messages Y Y NA Content, User Id, ts
Save/Favourite Y Y Y Content Id, ts
Share(In-app) Y Y NA Content Id, ts
Share(Across-app) Y N N Content Id, ts
Interests/Topics Y Y N List of topics
Time spent $^{*} NN^{*}$ $N^{*}$ Duration/frequency
User’s content Media Y Y Y Media file/URL
Text details Y Y Y title
Location Y Y Y Some place identifiers
Date time Y Y Y ts
Device - Y $Y^{g}$ Device model, OS
Other user interactions $^{*} NN^{*}$ $N^{*}$ Likes, Comments
Connections Y Y Y Username, ts
Login history Y Y $Y^{g}$ IP, ts
Current devices Y Y $Y^{g}$ User agent
Current camera NA Y NA Version/type
Location Y Y $Y^{g}$ Place identifiers
Account changes - Y N Type, Old, New values, ts
Advertisements Ads viewed N N Y Content Id, ts
Personalization $N^{*}$} N N Reasons why the ad was shown
Access to your data N Y N Which and how (in store visit etc.)
Miscellaneous Off-platform Y Y N Platform, ts, activity
Link history - Y - Link, ts
Cookies - Y - -
TABLE III: Data Transparency: Overview of shared information (as of Dec, 2024). $N^{*}$} denotes the details are found in the app but not in the GDPR dump. $Y^{g}$} denotes details are found in google’s DDP, but not in YouTube’s DDP. (ts: timestamp).
3.2 User engagement
One of the most important categories of information in the DDPs is how a user engages with social media platforms. It may include the content that a user watches, searches, or engages with implicitly or explicitly.
Watch history: Watch history is an ordered list of all the content that a user has watched before requesting the data. DDPs of all three platforms contain watch history data of the data subject. However, as Table I shows the data provided across the three platforms are widely varying. Among these, two points are strikingly different – (a) while YouTube provides watch history for the entire lifetime of a user on the platform, Instagram provides data for at most the last two weeks and TikTok provides data for around six months before the data request. These widely varying retention periods for watch history are provided without clear explanations for the criteria used, raising questions about full adherence to Article 15(1)(d) which requires information on data storage periods or the criteria for determining them. (b) While TikTok and YouTube provide the watch history all at once, Instagram segregates it into three files based on the type of content, i.e., ads, posts, and videos watched.
Like history: Similar to watch history, like history is an ordered list of all the contents that a user has liked before requesting the data. Table II demonstrates that there is a clear lack of consistency in how platforms currently implement Article 15(3) GDPR. The most surprising observation is YouTube does not even include like history in its DDP, despite ‘likes’ clearly constituting personal data processed by the platform and thus potentially falling under the scope of data subjects’ Right of Access. We would like to note that although Like history is not there in YouTube DDP, it is included in the Google DDP.
Time spent on platform: Despite showing such information on their apps, none of the platforms share the time users spend on platforms on a daily basis in the DDPs. Only TikTok has recently added a new field in the data, namely ‘Activity Summary’, which enlists the number of videos a user has commented on, shared, or watched until the end.
Other usage activities on platform:  Table III shows other usage activities. Most of the activities, e.g., comment, search, save, share, and writing a message are recorded by all platforms where the features are applicable. However, sharing across applications, i.e., when a user shares the video on other social media platforms or copies the link for posting elsewhere, is only shared in TikTok DDPs. Also, while both TikTok and Instagram maintain a list of inferred interests,YouTube DDPs do not have them.
3.3 User’s content